A block of mass 2m is intially at rest on a rough inclined plane where ? < 30 and is connected to an object with mass m as shown. the rope may be considerd massless, and the pulley may be considerd frictionless . the cofficient of static friction between the block and the plabe is us and the cofficient of kinetic friction between the block and the plane is uk.

A block of mass 2m is intially at rest on a rough

1. What is the magnitude of the static frictional force acting on the block resting on the slope?

A) mgsin ? B) usmgcos ? C) mg( 1 - us sin ? ) D) mg(1- us cos ? ) E) mg(1- 2sin ? )

2. What is the magnitude of the tension in the rope jpining the masses ?

A) mgsin ?   B) usmgcos ?   C) mg( 1 - sin ? ) D) mg(1- us cos ? ) E) mg

3. If the rope were cut between theblock and the pulley , and the block on the inclined plane began to slide down, what would be the magnitude of the acceleration of this block down the plane?

A) g B) g( sin ? - uksin ? )   C) g - 2 ukcos ? D) 2g (tan ? - uk ) E) 2g (sin ? - ukcos ?)
Answer #1

1)


equation of motion for mass "m" is,


tension T=m*g


equation of motion for mass "2m" is,


fs=T-2*mg*sin(theta)


fs=T-2*mg*sin(theta)


fs=T-2*mg*sin(theta)


fs=m*g*(1-2*sin(theta)) ------> option (E) is correct


2)


magnitude of the tension in the rope is, T=m*g ---->


option (E) is correct


3)

2m*a=2m*g*sin(theta)-uk*N


2m*a=2m*g*sin(theta)-uk*(2*mg*cos(theta))


a=g*(sin(theta)-uk*cos(theta))


acceleration is a=g*(sin(theta)-uk*cos(theta)) ---->